OpenStudy (anonymous):

Mr. Barnes writes the pattern for his math class on the board. 0, 2, 8, 18, 32, ... The class determines that each term of the pattern is twice a perfect square. Therefore, each term of the pattern is 2(n – 1)2, where n represents the term number. Which term is equal to 800? the 19th term the 20th term the 21st term the 22nd term

OpenStudy (anonymous):

it's 2(n-1)^2, right? not times 2 at the end?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

Alright, first off, set the equation equal to 800 2(n-1)^2 = 800 (n-1)^2 = 400 n-1 = squareroot(400) n = squareroot(400) - 1

OpenStudy (anonymous):

I think that's 19

OpenStudy (anonymous):

The rest of them you just have to plug into the calculator and they work.